OpenStudy (howard-wolowitz):

In a recent survey of 50 ten year olds, it was discovered that 14 liked both McDonalds and Burger King. Overall, 36 of the children surveyed liked McDonalds and 24 liked Burger King. How many children surveyed did not like niether restaurant? So statistical thinking to support answer!

hartnn (hartnn):

what have you tried?

hartnn (hartnn):

\((A' \cap B' ) = (A \cup B)' \\ A\cup B = A+B-A \cap B\)

OpenStudy (howard-wolowitz):

well this: N(not McD and not BK)=Total N - N(McD or BK) = 50-46 = 4

OpenStudy (howard-wolowitz):

@hartnn i did work son

hartnn (hartnn):

ok, 50 - (36+24-14) = 50 - 46 = 4 yep, you're correct :)